OpenStudy (anonymous):

what is the result of isolating y^2 in the equation (x+4)^2+y^2=22

OpenStudy (anonymous):

y^2=-x^2+6

OpenStudy (anonymous):

simply move (x+4)^2 to the other side by subtracting it from both sides

OpenStudy (anonymous):

if you wanted to isolate just 'y' the answer would be y=-x+7

OpenStudy (anonymous):

these are the obtions A. y^2=-x^2-8x+16 B.y^2=22-x^2 C.y^2=x^2-8x+6 D.y^2=-x^2-8x+6 @mrskipped8thgrade

OpenStudy (anonymous):

oh my bad. (x+4)(x+4)=x^2+8x+16

OpenStudy (anonymous):

so the answer is y^2=-x^2-8x+6

OpenStudy (anonymous):

D. is the correct answer

OpenStudy (anonymous):

thankssss !

OpenStudy (anonymous):

no problem, anytime.